# references.bib

### v2

@@ Abstract
Image-event joint depth estimation methods leverage complementary modalities for robust perception, yet face challenges in generalizability stemming from two factors: 1) limited annotated image-event-depth datasets causing insufficient cross-modal supervision, and 2) inherent frequency mismatches between static images and dynamic event streams with distinct spatiotemporal patterns, leading to ineffective feature fusion. To address this dual challenge, we propose Frequency-decoupled Unified Self-supervised Encoder (FUSE) with two synergistic components: The Parameter-efficient Self-supervised Transfer (PST) establishes cross-modal knowledge transfer through latent space alignment with image foundation models, effectively mitigating data scarcity by enabling joint encoding without depth ground truth. Complementing this, we propose the Frequency-Decoupled Fusion module (FreDFuse) to explicitly decouple high-frequency edge features from low-frequency structural components, resolving modality-specific frequency mismatches through physics-aware fusion. This combined approach enables FUSE to construct a universal image-event encoder that only requires lightweight decoder adaptation for target datasets. Extensive experiments demonstrate state-of-the-art performance with $14\%$ and $24.9\%$ improvements in Abs.Rel on MVSEC and DENSE datasets. The framework exhibits remarkable zero-shot adaptability to challenging scenarios including extreme lighting and motion blur, significantly advancing real-world deployment capabilities. The source code for our method is publicly available at: https://github.com/sunpihai-up/FUSE.
@@ I Introduction
Modern monocular depth estimation (MDE) increasingly utilizes multi-modal sensing to address the limitations of single-sensor systems [ramnet, xu2024unveiling], particularly by integrating conventional image sensors with emerging event cameras. Traditional image sensors, which operate using fixed-exposure light integration, effectively preserve structural information but suppress high-frequency temporal variations. This characteristic makes them suitable for static scene perception but limits their performance in dynamic scenarios. In contrast, event cameras asynchronously detect pixel-wise brightness changes with microsecond resolution [zhu2018ev], making them highly effective for capturing high-frequency motion but incapable of detecting static structures with subthreshold intensity gradients. The inherent complementarity in the frequency domain between these two modalities highlights the potential of image-event fusion as a promising approach for achieving robust depth estimation in dynamic environments.
Figure 1: Demonstration of FUSE’s superior performance in challenging conditions. (a) and (c) show the input image and event data, with blue/red in (c) indicating brightness decrease/increase. When image data fails due to low light and blur, event data provides complementary dynamic information. Our method (d) leverages multimodal synergy, recovering the traffic light missed by the state-of-the-art image depth model [depthanything2], as highlighted in the orange box.
@@
Recent studies [ramnet, srfnet, pcdepth] have explored leveraging the complementary characteristics of image and event modalities for joint MDE. These methods typically encode both modalities into a unified feature space, followed by customized feature fusion modules for depth prediction. However, their reliance on supervised learning with depth ground truth from image-event pairs has revealed poor generalization capabilities. This limitation arises from the scarcity and limited diversity of existing labeled image-event datasets [mvsec] compared to large-scale image-depth datasets such as NYUv2 [nyuv2] and KITTI [kitti].
Additionally, although the frequency-domain complementarity between images and events presents theoretical advantages, mismatches in their frequency characteristics introduce significant challenges for feature fusion. Conventional fusion strategies risk destructive interference [chen2024frequency]: high-frequency event features can disrupt the structural continuity of image representations, while low-frequency image components may suppress critical motion cues captured by events. This inherent frequency divergence, combined with limited training data, poses a dual challenge—requiring models to bridge the modality gap with insufficient supervision and resolve conflicting frequency signatures during feature integration.
Inspired by the impressive generalization capabilities of image-based depth foundation models (e.g., Depth Anything [depthanything1]), which are trained on millions of images, we propose a potential solution: transferring their well-learned image-depth priors to bridge the gap between event-based and depth-based representations. This insight motivates our two-fold innovation: 1) addressing data scarcity through parameter-efficient transfer learning, and 2) resolving intrinsic frequency mismatches via frequency decoupling mechanisms.
@@
To validate the effectiveness of FUSE, we conduct extensive experiments on the MVSEC and DENSE datasets, demonstrating state-of-the-art (SOTA) performance in MDE. As shown in Fig. 1, FUSE outperforms the current SOTA image depth model [depthanything2] under zero-shot settings, particularly in night-time driving scenarios affected by motion blur, highlighting its robustness. In summary, our contributions are as follows:
@@ II-A Image-based Monocular Depth Estimation
Deep learning-based methods have become the dominant paradigm in MDE. Eigen et al. [eigen] were the first to introduce a multi-scale fusion network for depth prediction through regression. Subsequent studies have improved depth estimation accuracy by reformulating the regression task as classification [adabins, binsformer, shao2023iebins], introducing novel network architectures [dpt], and incorporating additional prior knowledge [nddepth, newcrf]. Despite these advancements, the generalization capability of such models remains a major challenge.To address this limitation, recent research [midas, metric3d] has focused on training models on large-scale datasets aggregated from diverse sources, leading to substantial improvements. A milestone in this direction is MiDaS [midas], which mitigates the impact of varying depth scales across datasets by employing an affine-invariant loss. More recently, the DepthAnything series [depthanything1, depthanything2] has further enhanced the utilization of unlabeled data alongside large-scale labeled datasets, achieving remarkable performance. However, due to the inherent limitations of traditional image cameras, image-based MDE methods are still susceptible to factors such as lighting conditions and motion blur, which can degrade their performance [imgbad1, imgbad2, imgbad3].
@@ II-B Event-based Monocular Depth Estimation
Event-based MDE methods have attracted attention for their effectiveness under extreme lighting and high-speed motion conditions, making them suitable for applications in autonomous driving and robotic navigation [eventsurvey]. E2Depth [e2depth] was the first method to propose dense MDE using event data, employing a recurrent encoder-decoder architecture. Mixed-EF2DNet [mixed] introduced a flow network to capture temporal information effectively, while EReFormer [ereformer] leveraged a Transformer-based architecture for improved performance. Additionally, HMNet [hmnet] incorporated multi-level memory units to handle long-term dependencies efficiently. Nonetheless, these methods continue to face challenges due to the sparsity of event data and the limited availability of training datasets [eventbad].
@@ II-C Image-Event Fusion
The complementary characteristics of event and image data have been leveraged to enhance various visual tasks, including semantic segmentation [zhang2023cmx, zhang2023cmx], object detection [zhang2021object, zhou2023rgb], and depth estimation [ramnet, srfnet]. RAMNet [ramnet] employs an RNN-based approach to effectively utilize asynchronous event data, while EVEN [shi2023even] improves low-light performance by fusing modalities prior to inputting them into an MDE network. SRFNet [srfnet] guides modality interaction using spatial reliability masks, and PCDepth [pcdepth] integrates features at the modality level for more effective fusion. Our FUSE is the first to leverage knowledge transfer from an image foundation model for both feature encoding and fusion, significantly enhancing robustness and accuracy in image-event joint depth estimation.
@@ III-A Overview
TABLE I: Quantitative results on the MVSEC datast [mvsec]. $\downarrow$ indicates lower is better and denotes higher is better. The best results are highlighted in bold, while the second-best outcomes are underlined.
@@ III-B Event Representation
The event stream is first transformed into a voxel grid representation [ramnet] to make it compatible with existing network architectures. Each event $e_{n}$ is represented by the tuple $(x_{n},y_{n},t_{n},p_{n})$, which means that at time $t_{n}$, the log intensity change at the pixel position $(x_{n},y_{n})$ exceeds a threshold. The polarity $p_{n}\ (\pm 1)$ indicates whether the intensity increases or decreases. As such, the asynchronous event stream $\mathbf{S}=\{e_{n}\}_{n=0}^{N-1}$ must be transformed into a frame-like representation to be compatible with existing network architectures. Following prior works [ramnet], [srfnet], [pcdepth], [e2depth] we convert the temporal stream of events into a voxel grid [gehrig2019end].
@@
$\mathit{V}(x,y,t)=\sum p_{i}\delta\left(x-x_{i},y-y_{i}\right)\max\left\{0,1-\left|t-t_{i}^{*}\right|\right\}$ (1)
@@ III-C Parameter-efficient Self-supervised Transfer
Stage I: Parameter-Efficient Feature Alignment: In this stage, we align event data representations with the image domain’s latent space. We initialize $\mathcal{E}_{E}$ and $\mathcal{D}_{E}$ using weights from Depth Anything V2 [depthanything2] to inherit its geometric understanding ability. To enable efficient adaptation while preventing catastrophic forgetting [forgetcatastrophic], we implement LoRA[lora], where only the LoRA matrices and PatchEmbed Layer parameters (1.8% of $\mathcal{E}_{E}$) are updated. This allows adaptation to event data while retaining the foundation model’s generalization capability. The alignment is driven by a compound objective function (detailed in Sec. III-E) that enforces consistency in both output space (depth prediction) and latent representations.
Stage II: Robust Feature Fusion: In this stage, we focus on training the FreDFuse (Sec. III-D) in the image-event joint estimator to robustly fuse multi-modal features from image and event. The image encoder $\mathcal{E}_{I}$ and decoder $\mathcal{D}$ are inherited from the image foundation model, e.g., Depth Anything V2, while the event encoder $\mathcal{E}_{E}$ is obtained from Stage I. Only the FreDFuse is trainable during this stage. To improve the robustness and generalization of the proposed method, we also generate randomly degraded image-event pairs as inputs during training. The degraded inputs incorporate diverse noise types, including brightness adjustment, overexposure, motion blur, and occlusion masking. This systematic degradation strategy compels the feature fusion module to identify robust cross-modal relationships that withstand various sensor-specific degradations, thereby enhancing the model’s generalizability and resilience. Here, we use the same objective as Eq. 10 to maintain cross-modal consistency.
@@ III-E Optimization Objectives
where ${L}_{1}=||{d}-{d}^{*}||_{1}$ aligns output predictions, and ${L}_{\cos}=\left[1-\cos(\mathbf{F},\mathbf{F}^{*})\right]\cdot\mathbb{I}({\alpha\leq\cos(\mathbf{F},\mathbf{F}^{*})\leq\beta})$ aligns latent representations. Here, $\alpha=0.2$, $\beta=0.85$, and $\mathbb{I}(\cdot)$ ensure loss activation within the specified similarity range, preventing over-regularization.
Optimization Objective of Target Datasets: When transferring to target datasets with metric depth labels, we replace pseudo-supervision with the Scale-invariant Logarithmic (SiLog) loss [eigen]. Given a predicted depth map $\mathbf{d^{*}}$ and depth ground truth $\mathbf{d}$, the loss is computed over the pixels as follows:
@@
Figure 4: Qualitative analysis of the MVSEC dataset, outdoor_day1 scene. (a) and (b) show the input image and event data; (c) depicts the image-event joint estimation by HMNet[hmnet]; (d) shows the event-only estimation by HMNet; (e) presents our proposed FUSE; (f) shows the depth ground truth.
Figure 5: Qualitative analysis of the MVSEC dataset, outdoor_night1 scene. (a) and (b) show the input image and event data; (c) depicts the image-event joint estimation by HMNet[hmnet]; (d) shows the event-only estimation by HMNet; (e) presents our proposed FUSE; (f) shows the depth ground truth.
Figure 6: Qualitative comparison of zero-shot predictions when the image modality is severely degraded.(a) denotes the original image, (b) represents the image degraded by low light, overexposure, and occlusion masks, (c) denotes the event data, and (d) and (e) show the predicted results from DepthAnything V2 [depthanything2] and our FUSE, respectively.
@@ IV Experiment
To validate the effectiveness of our proposed FUSE framework in addressing the scarcity of event-depth data and modality disparity, we conduct comprehensive experiments under strictly controlled conditions. Our evaluation protocol emphasizes three critical aspects: (1) Generalization capability across real-world (MVSEC) and synthetic (DENSE) scenarios; (2) Performance constraints with frozen image-event joint encoders; (3) Computational efficiency in multi-modal fusion.
@@ IV-A Experimental Settings
We utilize EventScape [ramnet] for image-event pairs in PST (without depth ground truth). We evaluate on MVSEC [mvsec] and DENSE [e2depth] datasets, where MVSEC is a real-world scenario while EventScape and DENSE are synthetic scenarios.
MVSEC Dataset: The MVSEC dataset uses a pair of DAVIS cameras to capture grayscale images and event data at a resolution of $346\times 260$. Depth ground truth is recorded using a LiDAR sensor at 20 Hz. Grayscale images are captured at 10 Hz during the day and 45 Hz at night.
Metrics: Following [pcdepth], we evaluate performance using the metrics: Absolute Relative Error (Abs.Rel), Root Mean Square Error (RMSE), Logarithmic Squared Error (RMSELog), average absolute depth errors at different cut-off depth distances (i.e., 10m, 20m, and 30m), and accuracy ($\delta<1.25^{n},n=1,2,3$).
Implementation Details: The model is trained using the Adam optimizer with a learning rate of 5e-5, implemented in PyTorch on two RTX 3090ti GPUs with a batch size of 48. PST on EventScape [ramnet] is trained for 10 epochs, and fine-tuning on MVSEC and DENSE for 20 epochs. Manual alignment is needed for the MVSEC dataset due to the asynchronous nature of images, events, and depth labels. As in [pcdepth], event data from the first 50 ms prior to each depth ground truth is paired with the most recent image to form the data pairs.
@@ IV-B Evaluation on MVSEC and DENSE Datasets
We report quantitative results on the real-world MVSEC [mvsec] and synthetic DENSE [e2depth] datasets. For MVSEC, we use the outdoor_day2, outdoor_night2, and outdoor_night3 sequences for training, and outdoor_day1 and outdoor_night1 for testing, with depth range from 1.97 to 80 meters [pcdepth]. For DENSE, we use the provided training and test sets, with depth range from 3.34 to 1000 meters [ramnet, srfnet]. We freeze the image-event joint encoder and train only the depth decoder.
Analysis on MVSEC Dataset: Tab. I presents a quantitative comparison between our method and SOTA methods for event-based estimation and image-event joint estimation on the MVSEC dataset. In the outdoor_day1 scene, our method outperforms the best image-event joint estimation approach [pcdepth] by $14\%$ and $10.2\%$ in Abs.Rel and RMSELog, respectively. The progress achieved by freezing the encoder and dealing with the significant disparity between source and target data clearly demonstrates the effectiveness of our proposed FUSE in addressing the scarcity of event depth data and providing robust and accurate depth estimation. Fig. 4 and Fig. 5 present qualitative comparisons with HMNet[hmnet] for the outdoor_day1 and outdoor_night1 scenes, respectively. FUSE provides finer and more stable depth predictions. Compared to HMNet, our method demonstrates a superior ability to preserve structural details, delivering more consistent and accurate predictions on both buildings and vegetation.
Analysis on DENSE Dataset: Tab. II presents the quantitative results on the synthetic DENSE dataset [e2depth]. Our method achieves improvements of $24.9\%$ in Abs.Rel and $33.4\%$ in RMSELog over previous approaches. For the average absolute depth error at the truncated distance, our method ranks either the best or the second-best among the evaluated methods. Fig. 6 shows a qualitative comparison under extreme damage to the image modality. Our FUSE is still able to provide stable depth estimation even when some regions of the image are completely destroyed.
TABLE II: Comparison of performances on synthetic DENSE dataset [e2depth]

### v3

# FUSE: Label-Free Image-Event Joint Monocular Depth Estimation via Frequency-Decoupled Alignment and Degradation-Robust Fusion

@@ Abstract
Image-event joint depth estimation methods leverage complementary modalities for robust perception, yet face challenges in generalizability stemming from two factors: 1) limited annotated image-event-depth datasets causing insufficient cross-modal supervision, and 2) inherent frequency mismatches between static images and dynamic event streams with distinct spatiotemporal patterns, leading to ineffective feature fusion. To address this dual challenge, we propose Frequency-decoupled Unified Self-supervised Encoder (FUSE) with two synergistic components:
The Parameter-efficient Self-supervised Transfer (PST) leverages image foundation models for cross-modal knowledge transfer, effectively mitigating data scarcity by enabling joint encoding without depth ground truth.
Complementing this, the Frequency-Decoupled Fusion module (FreDFuse) resolves modality-specific frequency mismatches by decoupling features into high- and low-frequency bands and then performing a guided cross-attention fusion, where the modality dominant in each band steers the integration. This combined approach enables FUSE to construct a universal image-event encoder that only requires lightweight decoder adaptation for target datasets. Extensive experiments demonstrate state-of-the-art performance with $14\%$ and $24.9\%$ improvements in Abs.Rel on MVSEC and DENSE datasets.
The framework exhibits remarkable robustness and generalization in challenging scenarios, including extreme lighting and motion blur, significantly advancing its real-world deployment capabilities. The source code for our method is publicly available at: https://github.com/sunpihai-up/FUSE.
@@ I Introduction
Modern monocular depth estimation (MDE) increasingly utilizes multi-modal sensing to address the limitations of single-sensor systems [6, 24], particularly by integrating conventional image sensors with emerging event cameras. Traditional image sensors, which operate using fixed-exposure light integration, effectively preserve structural information but suppress high-frequency temporal variations. This characteristic makes them suitable for static scene perception but limits their performance in dynamic scenarios. In contrast, event cameras asynchronously detect pixel-wise brightness changes with microsecond resolution [32], making them highly effective for capturing high-frequency motion but incapable of detecting static structures with subthreshold intensity gradients. The inherent complementarity in the frequency domain between these two modalities highlights the potential of image-event fusion as a promising approach for achieving robust depth estimation in dynamic environments.
Figure 1: Demonstration of FUSE’s superior performance in challenging conditions. (a) and (c) show the input image and event data, with blue/red in (c) indicating brightness decrease/increase. When image data fails due to low light and blur, event data provides complementary dynamic information. Our method (d) leverages multimodal synergy, recovering the traffic light missed by the state-of-the-art image depth model [25], as highlighted in the orange box.
@@
Recent studies [6, 16, 13] have explored leveraging the complementary characteristics of image and event modalities for joint MDE. These methods typically encode both modalities into a unified feature space, followed by customized feature fusion modules for depth prediction. However, their reliance on supervised learning with depth ground truth from image-event pairs has revealed poor generalization capabilities. This limitation arises from the scarcity and limited diversity of existing labeled image-event datasets [33] compared to large-scale image-depth datasets such as NYUv2 [23] and KITTI [8].
Additionally, although the frequency-domain complementarity between images and events presents theoretical advantages, mismatches in their frequency characteristics introduce significant challenges for feature fusion. Conventional fusion strategies risk destructive interference [1]: high-frequency event features can disrupt the structural continuity of image representations, while low-frequency image components may suppress critical motion cues captured by events. This inherent frequency divergence, combined with limited training data, poses a dual challenge—requiring models to bridge the modality gap with insufficient supervision and resolve conflicting frequency signatures during feature integration.
Inspired by the impressive generalization capabilities of image-based depth foundation models (e.g., Depth Anything [26]), which are trained on millions of images, we propose a potential solution: transferring their well-learned image-depth priors to bridge the gap between event-based and depth-based representations. This insight motivates our two-fold innovation: 1) addressing data scarcity through parameter-efficient transfer learning, and 2) resolving intrinsic frequency mismatches via frequency decoupling mechanisms.
@@
To validate the effectiveness of FUSE, we conduct extensive experiments on the MVSEC and DENSE datasets, demonstrating state-of-the-art (SOTA) performance in MDE. As shown in Fig. 1, FUSE outperforms the current SOTA image depth model [25] under zero-shot settings, particularly in night-time driving scenarios affected by motion blur, highlighting its robustness. In summary, our contributions are as follows:
@@ II-A Image-based Monocular Depth Estimation
Deep learning-based methods have become the dominant paradigm in MDE. Eigen et al. [3] were the first to introduce a multi-scale fusion network for depth prediction through regression. Subsequent studies have improved depth estimation accuracy by reformulating the regression task as classification [4, 12], introducing novel network architectures [17], and incorporating additional prior knowledge [20, 29]. Despite these advancements, the generalization capability of such models remains a major challenge.To address this limitation, recent research [18] has focused on training models on large-scale datasets aggregated from diverse sources, leading to substantial improvements. A milestone in this direction is MiDaS [18], which mitigates the impact of varying depth scales across datasets by employing an affine-invariant loss. More recently, the DepthAnything series [26, 25] has further enhanced the utilization of unlabeled data alongside large-scale labeled datasets, achieving remarkable performance. Concurrently, other research [27] has focused on improving generalization through model-centric approaches, such as employing stabilized adversarial training to boost robustness against domain shifts. However, due to the inherent limitations of traditional image cameras, image-based MDE methods are still susceptible to factors such as lighting conditions and motion blur, which can degrade their performance [2].
@@ II-B Event-based Monocular Depth Estimation
Event-based MDE methods have attracted attention for their effectiveness under extreme lighting and high-speed motion conditions, making them suitable for applications in autonomous driving and robotic navigation [5]. E2Depth [10] was the first method to propose dense MDE using event data, employing a recurrent encoder-decoder architecture. Mixed-EF2DNet [21] introduced a flow network to capture temporal information effectively, while EReFormer [14] leveraged a Transformer-based architecture for improved performance. Additionally, HMNet [9] incorporated multi-level memory units to handle long-term dependencies efficiently. Nonetheless, these methods continue to face challenges due to the sparsity of event data and the limited availability of training datasets [19].
@@ II-C Image-Event Fusion
The complementary characteristics of event and image data have been leveraged to enhance various visual tasks, including semantic segmentation [30, 28], object detection [31], and depth estimation [6, 16]. RAMNet [6] employs an RNN-based approach to effectively utilize asynchronous event data, while EVEN [22] improves low-light performance by fusing modalities prior to inputting them into an MDE network. SRFNet [16] guides modality interaction using spatial reliability masks, and PCDepth [13] integrates features at the modality level for more effective fusion. Our FUSE is the first to leverage knowledge transfer from an image foundation model for both feature encoding and fusion, significantly enhancing robustness and accuracy in image-event joint depth estimation.
@@ III-A Overview
TABLE I: Quantitative results on the MVSEC datast [33]. $\downarrow$ indicates lower is better and denotes higher is better. The best results are highlighted in bold, while the second-best outcomes are underlined.
@@ III-B Event Representation
The event stream is first transformed into a voxel grid representation [6] to make it compatible with existing network architectures. Each event $e_{n}$ is represented by the tuple $(x_{n},y_{n},t_{n},p_{n})$, which means that at time $t_{n}$, the log intensity change at the pixel position $(x_{n},y_{n})$ exceeds a threshold. The polarity $p_{n}\ (\pm 1)$ indicates whether the intensity increases or decreases. As such, the asynchronous event stream $\mathbf{S}=\{e_{n}\}_{n=0}^{N-1}$ must be transformed into a frame-like representation to be compatible with existing network architectures. Following prior works [6], [16], [13], [10] we convert the temporal stream of events into a voxel grid [7].
@@
$\mathit{V}(x,y,t)=\sum p_{i}\delta\left(x-x_{i},y-y_{i}\right)\max\left\{0,1-\left|t-t_{i}^{*}\right|\right\},$ (1)
@@ III-C Parameter-efficient Self-supervised Transfer
Stage I: Parameter-Efficient Feature Alignment: In this stage, we align event data representations with the image domain’s latent space. We initialize $\mathcal{E}_{E}$ and $\mathcal{D}_{E}$ using weights from Depth Anything V2 [25] to inherit its geometric understanding ability. To enable efficient adaptation while preventing catastrophic forgetting [15], we implement LoRA[11], where only the LoRA matrices and PatchEmbed Layer parameters (1.8% of $\mathcal{E}_{E}$) are updated. This allows adaptation to event data while retaining the foundation model’s generalization capability. The alignment is driven by a compound objective function (detailed in Sec. III-E) that enforces consistency in both output space (depth prediction) and latent representations.
Stage II: Robust Feature Fusion: This stage focuses on training only the FreDFuse module (Sec. III-D) for robust multi-modal fusion, while keeping other components frozen: the image encoder ($\mathcal{E}_{I}$) and decoder ($\mathcal{D}$) from a foundation model (e.g., Depth Anything V2), and the event encoder ($\mathcal{E}_{E}$) from Stage I. To improve robustness and generalization, we employ a systematic data augmentation strategy by training on randomly degraded image-event pairs. This includes global brightness adjustments (factor $\in[0.5,1.5]$) and localized degradations applied to 50% of the pairs in random $20\%\times 20\%$ regions (detailed in Table II). This strategy forces the fusion module to learn resilient cross-modal correlations, using the objective from Eq. 10 to enforce consistency.
TABLE II: Parameterization of the random degradation strategy.
  Degradation Type Parameters Applicable Modality Gaussian Blur Kernel size: $25\times 25$, $\sigma=2$ Image & Event Overexposure $\alpha=2.5$, $\beta=200$ Image only Occlusion Fill region with black pixels Image & Event
@@ III-E Optimization Objectives
[…]
Optimization Objective of Target Datasets: When transferring to target datasets with metric depth labels, we replace pseudo-supervision with the Scale-invariant Logarithmic (SiLog) loss [3]. Given a predicted depth map $\mathbf{d^{*}}$ and depth ground truth $\mathbf{d}$, the loss is computed over the pixels as follows:
@@
Figure 4: Qualitative analysis of the MVSEC dataset, outdoor_day1 scene. (a) and (b) show the input image and event data; (c) depicts the image-event joint estimation by HMNet[9]; (d) shows the event-only estimation by HMNet; (e) presents our proposed FUSE; (f) shows the depth ground truth.
Figure 5: Qualitative analysis of the MVSEC dataset, outdoor_night1 scene. (a) and (b) show the input image and event data; (c) depicts the image-event joint estimation by HMNet[9]; (d) shows the event-only estimation by HMNet; (e) presents our proposed FUSE; (f) shows the depth ground truth.
Figure 6: Qualitative comparison of zero-shot predictions when the image modality is severely degraded.(a) denotes the original image, (b) represents the image degraded by low light, overexposure, and occlusion masks, (c) denotes the event data, and (d) and (e) show the predicted results from DepthAnything V2 [25] and our FUSE, respectively.
@@ IV Experiment
To validate the effectiveness of our proposed FUSE framework in addressing the scarcity of event-depth data and modality disparity, we conduct comprehensive experiments under strictly controlled conditions.
@@ IV-A Experimental Settings
We utilize EventScape [6] for image-event pairs in PST (without depth ground truth). We evaluate on MVSEC [33] and DENSE [10] datasets, where MVSEC is a real-world scenario while EventScape and DENSE are synthetic scenarios.
Metrics: Following [13], we evaluate performance using the metrics: Absolute Relative Error (Abs.Rel), Root Mean Square Error (RMSE), Logarithmic Squared Error (RMSELog), average absolute depth errors at different cut-off depth distances (i.e., 10m, 20m, and 30m), and accuracy ($\delta<1.25^{n},n=1,2,3$).
Implementation Details: The model is trained using the Adam optimizer with a learning rate of 5e-5, implemented in PyTorch on two RTX 3090ti GPUs with a batch size of 48. PST on EventScape [6] is trained for 10 epochs, and fine-tuning on MVSEC and DENSE for 20 epochs. Manual alignment is needed for the MVSEC dataset due to the asynchronous nature of images, events, and depth labels. As in [13], event data from the first 50 ms prior to each depth ground truth is paired with the most recent image to form the data pairs.
@@ IV-B Evaluation on MVSEC and DENSE Datasets
We evaluate our method on the real-world MVSEC [33] and synthetic DENSE [10] datasets. For MVSEC, we train on outdoor_day2, outdoor_night2, and outdoor_night3, and test on outdoor_day1 and outdoor_night1 (depth range: 1.97-80m) [13]. For DENSE, we use its official splits (depth range: 3.34-1000m) [6, 16]. During training, we freeze the joint encoder and only train the depth decoder.
Analysis on MVSEC Dataset: Tab. I presents a quantitative comparison between our method and SOTA methods for event-based estimation and image-event joint estimation on the MVSEC dataset. In the outdoor_day1 scene, our method outperforms the best image-event joint estimation approach [13] by $14\%$ and $10.2\%$ in Abs.Rel and RMSELog, respectively. The progress achieved by freezing the encoder and dealing with the significant disparity between source and target data clearly demonstrates the effectiveness of our proposed FUSE in addressing the scarcity of event depth data and providing robust and accurate depth estimation. Fig. 4 and Fig. 5 present qualitative comparisons with HMNet[9] for the outdoor_day1 and outdoor_night1 scenes, respectively. FUSE provides finer and more stable depth predictions. Compared to HMNet, our method demonstrates a superior ability to preserve structural details, delivering more consistent and accurate predictions on both buildings and vegetation.
Analysis on DENSE Dataset: Tab. III presents the quantitative results on the synthetic DENSE dataset [10]. Our method achieves improvements of $24.9\%$ in Abs.Rel and $33.4\%$ in RMSELog over previous approaches. For the average absolute depth error at the truncated distance, our method ranks either the best or the second-best among the evaluated methods. Fig. 6 shows a qualitative comparison under extreme damage to the image modality. Our FUSE is still able to provide stable depth estimation even when some regions of the image are completely destroyed.
TABLE III: Comparisons on synthetic DENSE dataset [10]